OpenStudy (anonymous):

let f(x,y)= x^2cos(piy)-ln(2xy) find fx(2,3)+ fy(2,3)

OpenStudy (anonymous):

\[x^2 cos(\pi y)- ln(2xy)\] \[f_x= 2x cos(\pi y)-\frac{2y}{2 x y}\] \[f_x= 2x cos(\pi y)-\frac{1}{ x }\]

OpenStudy (anonymous):

\[f_y=-x^2\pi sin(\pi y)-\frac{1}{y} )\]

OpenStudy (anonymous):

\[f_x(2,3)=2(2)cos(\pi 3)- \frac{1}{2}\]

OpenStudy (anonymous):

i got -29/6 thanks so much

OpenStudy (anonymous):

enjoying calc 3?

OpenStudy (anonymous):

hahahah....fun stuff...enjoying tutoring?

OpenStudy (anonymous):

not really, not enough hours

OpenStudy (anonymous):

:( is it a parttime job? can u help me with another problem?

OpenStudy (anonymous):

post it, I will come back in about half an hour

OpenStudy (anonymous):

k thanks :)